in a right angled triangle the length of one of the sides forming the right angle is 4 cm more than twice the length of the other side. if the area of triangle is 120 cmsq. find its perimeter?
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

The legs of the triangle are x and 2x+4 cm.

(x)(2x+4) = 2 * 120
2x^2 + 4x - 240 = 0
x^2 + 2x - 120 = 0
(x+12)(x - 10) = 0

The sides are 10 cm and 24 cm, so the perimeter is 10+24+26 = 50cm.

Answer by ikleyn(52817)

.

Is the "other side" a leg, too ?



BY THE WAY

For right triangle,  the standard naming of its  sides forming the right angle  is one short English word of three letters,

and this word is  "leg".